### v1

@@ A.5 Comparison of FID and CLIP scores
Type Method Training Dataset FID Score CLIP Score
Single StableDiffusion [4] RSICD 103.4 26.1
StableDiffusion [4] DIOR 228.1 26.2
InstanceDiffusion [34] RSICD 138.6 24.7
Text2Earth [17] Git-10M 24.5* -
DiffusionSat [12] fMoW 15.8* 17.2*
GeoSynth [25] US Cities 12.3* / 171.1 30.3* / 25.0
SatSynth [14] iSAID, LoveDA, OEM - -
MMO-IG [16] DIOR 34.5* -
AeroGen [15] HRSC,DIOR 38.6* -
Multi CRS-Diff [13] RSICD, fMoW 50.7* / 107.0 20.3* / 23.6
InstanceDiffusion [34] RSICD, RSITMD, UCMerced 123.1 25.0
ControlNet [32] EarthSynth-180K 183.8 25.9
EarthSynth EarthSynth-180K 198.7 26.1
A.6 Few-Shot Detection Results
Table 8 shows that open-source methods such as ETS [54] consistently outperform close-source baselines across all shot settings. For example, ETS achieves 37.5 percent mAP in the 10-shot setting, surpassing Detic at 15.4 percent and DE-ViT at 25.6 percent. Introducing synthetic data leads to notable improvements, especially when combined with real data. In the 5-shot setting, ETS improves from 29.3 percent using only real data to 33.9 percent with ControlNet-generated data and 34.1 percent with EarthSynth-generated data. EarthSynth consistently yields better results than ControlNet, suggesting higher-quality synthesis. These results highlight the effectiveness of augmenting real data with high-quality synthetic data for few-shot detection.
Table 8: The 1/5/10-shot detection results on the DIOR dataset.
Method Setting Data Usage mAP
1-shot Detic ViT-L [55] Close-Source Real 4.1
DE-ViT ViT-L [56] Close-Source Real 14.7
CD-ViTO [57] Close-Source Real 17.8
GroundingDINO Swin-B [45] Open-Source Real 11.7
ETS Swin-B [54] Open-Source Real 12.7
(ControlNet) ETS Swin-B [54] Open-Source Synth 9.2
(ControlNet) ETS Swin-B [54] Open-Source Real + Synth 13.1
(EarthSynth) ETS Swin-B [54] Open-Source Synth 9.5
(EarthSynth) ETS Swin-B [54] Open-Source Real + Synth 13.9
5-shot Detic ViT-L [55] Close-Source Real 12.1
DE-ViT ViT-L [56] Close-Source Real 23.4
CD-ViTO [57] Close-Source Real 26.9
GroundingDINO Swin-B [45] Open-Source Real 27.7
ETS Swin-B [54] Open-Source Real 29.3
(ControlNet) ETS Swin-B [54] Open-Source Synth 23.0
(ControlNet) ETS Swin-B [54] Open-Source Real + Synth 33.9
(EarthSynth) ETS Swin-B [54] Open-Source Synth 23.6
(EarthSynth) ETS Swin-B [54] Open-Source Real + Synth 34.1
10-shot Detic ViT-L [55] Close-Source Real 15.4
DE-ViT ViT-L [56] Close-Source Real 25.6
CD-ViTO [57] Close-Source Real 30.8
GroundingDINO Swin-B [45] Open-Source Real 36.4
ETS Swin-B [54] Open-Source Real 37.5
(ControlNet) ETS Swin-B [54] Open-Source Synth 25.3
(ControlNet) ETS Swin-B [54] Open-Source Real + Synth 40.2
(EarthSynth) ETS Swin-B [54] Open-Source Synth 26.4
(EarthSynth) ETS Swin-B [54] Open-Source Real + Synth 40.7
A.7 Examples of Generated Data
Figure 11 and Figure 12 present examples of the generated semantic segmentation and object detection data produced by the proposed method. From a visual perspective, segmentation labels are generally of higher quality compared to object detection labels, as the latter often rely on post-processing algorithms such as edge detection, which may introduce inaccuracies.
Figure 11: Some examples of generated semantic segmentation data.
@@ A.7 Examples of Generated Data
Figure 12: Some examples of generated object detection data.
A.8 Comparison of Generated Images
CRS-Diff. Due to its limited training data, CRS-Diff produces lower-quality images compared to models trained on larger and more diverse datasets.
GeoSynth. GeoSynth is designed for high-resolution image generation based on data from US cities, but its capacity to generate a wide range of categories is limited.
Stable Diffusion. Stable Diffusion lacks the ability to generate layout-controllable images, which limits its applicability in tasks that require precise spatial arrangement of objects, such as semantic segmentation or object detection data synthesis.
ControlNet and EarthSynth. ControlNet and EarthSynth enable the generation of images with explicit layout control and visually salient targets, making them well-suited for structured data synthesis tasks.
Figure 13: Comparison of generated images with different methods.
A.9 Guidance Scale Analysis
@@
Figure 14 shows how the guidance scale affects the CLIP score across two categories. As the scale increases, the CLIP Score generally rises at first, then levels off or slightly decreases. This suggests that moderate guidance scales lead to better alignment between the generated image and the text prompt. In the top examples, representing tennis courts, the CLIP Score reaches its peak around scale 4, with images showing clearer structure and improved object fidelity. In the bottom examples, representing playgrounds, the score is highest near scale 2 or 3, but have a poor image generated. Lower scales tend to produce blurry or semantically weak images, while higher scales enhance visual clarity but may reduce diversity. These results indicate that a moderate guidance scale, typically between 3 and 5, offers a good balance between semantic alignment and image quality. The results demonstrate that generation varies across different images, and tuning the guidance scale provides a simple way to control semantic accuracy and visual structure.

### v2

@@ A.2.3 Data Sources of EarthSynth-180K
A.3 Remote Sensing Diffusion Models
Table 8 compares remote sensing diffusion models across three prompt levels: semantic, region, and pixel. Most methods support only semantic-level prompts, offering global but coarse control. Some models introduce region-level prompts using bounding boxes to enhance spatial precision. Pixel-level prompts, such as masks and sketches, provide the most detailed control and are used by CRS-Diff, GeoSynth, and EarthSynth. EarthSynth uniquely combines semantic and pixel-level prompts, enabling high-level semantics and fine-grained spatial guidance. This reflects a shift toward more precise and controllable image generation.
High-Resolution Synthesis Text2Earth, DiffusionSat, and GeoSynth are representative methods to generate high-resolution satellite imagery. These models leverage diffusion-based generative frameworks or text-to-image architectures to reconstruct fine-grained spatial details, often guided by auxiliary inputs such as text descriptions, semantic maps, or multi-modal signals. Their primary applications lie in image restoration, super-resolution, cloud removal, and spectral enhancement, which are essential for improving satellite data’s visual and analytical quality in scientific and operational settings.
Method Visual Quality Semantic Richness Overall Score
ControlNet 79.51 23.76 51.64
EarthSynth (Ours)) 80.19 25.95 53.07
Table 10: Image scoring results on 100 diffusion-generated images using GPT-4.
Task-Oriented Synthesis CRS-Diff, SatSynth, AeroGen, and MMO-IG are designed with downstream utility, focusing on generating synthetic data tailored for specific tasks such as land cover classification, object detection, and change detection. These models incorporate task-specific priors, including class labels, semantic layouts, or instance-level masks, to guide the generation process. By aligning the synthesized data with the needs of target tasks, these methods enhance model generalization in low-resource scenarios, enable domain adaptation, and facilitate pretraining or fine-tuning of models in remote sensing applications.
A.4 More Experiments
A.4.1 Comparison of FID and CLIP scores
Table 9 compares fake data from different models using FID and CLIP scores. FID measures distribution distance to all images of the RSICD dataset, while CLIP evaluates semantic alignment. The results reveal a mismatch between FID and CLIP metrics across models. For example, GeoSynth shows low FID (12.3) and high CLIP (30.3), indicating strong visual and semantic quality. In contrast, StableDiffusion on RSICD or DIOR has a high FID (103.4, 228.1) but is similar to CLIP on EarthSynth (26.1). InstanceDiffusion improves FID with multi-source data, yet its CLIP score stays nearly the same (25.0 vs. 24.7), underscoring a gap between visual fidelity and semantic alignment. This suggests that FID alone is not reliable for assessing task-specific data quality.
A.4.2 Multi-modal LLM for Image Quality Evaluation
Image Scoring based on Multi-modal LLM. We use GPT-4 as a text-based tool to evaluate the generated images, since it can’t analyze pixels directly. We extract basic data like resolution, color mode, and estimated visual complexity based on color count, then turn this into a descriptive prompt that summarizes the image’s main features. Our prompt consists of two components: (1) a system prompt that directs GPT-4 as an expert evaluator capable of judging image quality and semantic richness based only on textual input. The system prompt states: “You are an expert image evaluator who assesses image quality and semantic richness from detailed descriptions alone. You do not require access to the actual image; provide a numeric score from 0 to 100 based solely on the description.” (2) a user prompt that supplies the image description and explicitly instructs GPT-4 to respond in the exact format: “Score: X. Reason: <brief explanation>”, where X is an integer between 0 and 100. The prompt explicitly forbids disclaimers such as “I am an AI and cannot view images.”
You are an expert image evaluator. Based ONLY on the description below,
rate the image on two aspects:
1. Visual Quality (integer score 0-100)
2. Semantic Richness (integer score 0-100)
Image description: <image description text>
Respond ONLY with exactly the following format and nothing else:
Visual Quality Score: <integer 0-100>
Semantic Richness Score: <integer 0-100>
Reason: <one brief sentence>
Image scoring is performed along two key dimensions: (1) visual quality, encompassing factors such as image clarity, resolution, and compositional coherence; and (2) semantic richness, including the quantity of meaningful elements, scene complexity, and the depth of emotional or narrative content. This structured prompt formulation promotes consistent and parsable outputs, reducing subjective variance. Numeric scores are extracted via regular expression matching, and average scores are computed over the entire image set to yield final metrics. The prompts are listed below:
Method CLIP Train Data Top1 Acc
Baseline (EarthSynth) - 49.07
+ RemoteCLIP RET-3 + DET-10 + SEG-4 48.13
+ CLIP Web 50.47
Table 11: CLIP-based scene classification performance on DOTA-v2 using different module configurations.
Experimental Setup and Results. We evaluated 100 generated images for each method. The assessment was performed using GPT-4, which rated the images along two dimensions: Visual Quality, reflecting the perceptual fidelity and aesthetic appeal, and Semantic Richness, indicating the depth and variety of semantic content present in the image. The Overall Average Score was computed as the arithmetic mean of the two individual scores.
As shown in Table 10, EarthSynth achieved a slightly higher visual quality score (80.19) than ControlNet (79.51), indicating marginally better perceptual quality. EarthSynth produced images with greater semantic richness, scoring 25.95 compared to 23.76 from ControlNet. Consequently, ControlNet achieved a higher overall average score of 53.07, compared to 51.64 for EarthSynth. These results suggest EarthSynth performs better at both visual quality and semantic richness. We believe its scoring results can only serve as a reference and are intended to provide one of the methods for diversified image quality evaluation.
Figure 10: EarthSynth over time-step training process.
A.4.3 Ablation Studies
Local Loss. We qualitatively analyze the role of local constraints by visualizing the training process of the EarthSynth model. As shown in Figure 11 and Figure 10, the loss function is defined as,
Setting $\gamma=10$ accelerates convergence and helps the model capture semantic mask information better. Compared to not using $\mathcal{L}_{\texttt{local}}$, incorporating it enables more effective layout control. We found it challenging to learn layout control without local constraints on the mask. For comparison fairness, ControlNet with $\mathcal{L}_{\texttt{local}}$ is used in object detection and semantic segmentation.
Figure 11: EarthSynth over time-step training process.
CF-Comp. As illustrated in Figure 11 and Figure 10, we qualitatively analyze the impact of local constraints by visualizing the training process of the EarthSynth model. A visual ablation study on ICS and TSS further highlights their complementary roles within our framework. Specifically, incorporating ICS encourages the model to generate more diverse and novel compositional combinations, enhancing creativity and variety. In contrast, applying TSS fosters the generation of semantically coherent and realistic compositions, improving overall plausibility. These findings also validate the effectiveness of our proposed EarthSynth with CF-Comp strategy, demonstrating its ability to balance novelty with semantic fidelity.
CLIP of R-Filter. Table 11 presents the Top-1 accuracy of different configurations of the EarthSynth model on the DOTA-v2 scene classification task. The baseline EarthSynth model achieves an accuracy of 49.07. When integrated with RemoteCLIP, performance slightly drops to 48.13, suggesting that RemoteCLIP may not effectively enhance EarthSynth in this context, possibly due to misalignment with remote sensing imagery. In contrast, incorporating the standard CLIP module leads to a significant improvement, reaching the highest accuracy of 50.47. This result highlights CLIP’s strong capability to boost scene classification performance. Also, this reacts to the semantic bias that RemoteCLIP carries in the remote sensing domain, which is not well used for filtering.
Figure 12: Examples of generated segmentation data.
A.5 More Visualization
A.5.1 Examples of EarthSynth-generated Data
Figure 12 and Figure 13 illustrate representative examples of the synthetic semantic segmentation and object detection datasets generated using the proposed EarthSynth framework. These visualizations show that the semantic segmentation labels exhibit higher precision and consistency across diverse land cover types. This can be attributed to the pixel-level supervision involved in the segmentation process, which enables more accurate delineation of object boundaries. Object detection annotations are less reliable because they rely on post-processing techniques like edge detection and bounding box generation. These methods often introduce noise or result in misalignment with the actual object extents, thereby reducing the overall annotation quality. This contrast highlights the relative robustness of the segmentation outputs and suggests that EarthSynth is particularly well-suited for applications where spatial accuracy and detailed contextual information are essential.
A.5.2 Comparison of Generated Images from Remote Sensing Diffusion Models
Figure 14 compares generated images from remote sensing diffusion models. The focus of the different methods for generating images is described below.
CRS-Diff. CRS-Diff exhibits relatively limited image quality due to its constrained training data. The restricted diversity and quantity of training samples negatively impact the model’s ability to generalize across various geographic regions and land cover types. As a result, the generated images often lack the visual fidelity and semantic richness observed in outputs from models trained on more comprehensive datasets.
GeoSynth. GeoSynth is tailored explicitly for generating high-resolution remote sensing images, with training data predominantly sourced from urban areas in the United States. While it excels in producing detailed imagery within this domain, its generative capacity is limited when synthesizing scenes outside this geographic or semantic scope. In particular, its ability to represent diverse land use categories or non-urban environments is relatively weak, restricting its applicability in global or multi-domain remote sensing tasks.
@@ A.5.2 Comparison of Generated Images from Remote Sensing Diffusion Models
Figure 13: Examples of generated object detection data. This includes the ability to generate data from different satellites.
Stable Diffusion. Although Stable Diffusion demonstrates strong general-purpose image generation capabilities, it lacks mechanisms for spatial layout control. This limitation is critical in remote sensing applications, where the accurate placement and arrangement of objects, such as buildings, roads, and vegetation, are essential for downstream tasks. The inability to control spatial semantics diminishes its utility in structured synthesis scenarios where geospatial coherence and object positioning matter.
ControlNet and EarthSynth. ControlNet and the proposed EarthSynth model support explicit spatial layout control, allowing for the guided generation of images with well-defined structures and localized semantic targets. This capability is particularly valuable in tasks such as data augmentation, simulation-based training, or synthetic dataset creation for segmentation and detection models. EarthSynth, in particular, further enhances visual realism while preserving layout fidelity, making it a powerful tool for generating structured, high-quality remote sensing imagery across diverse environments and object categories.
A.5.3 Guidance Scale Analysis
Figure 14: Comparison of different generation methods.
@@ A.5.3 Guidance Scale Analysis
Figure 15 shows how the guidance scale affects the CLIP score. As the scale increases, the CLIP Score rises first, then levels off or slightly decreases. This suggests that moderate guidance scales lead to better alignment between the generated image and the text prompt. In the top examples, representing tennis courts, the CLIP Score reaches its peak around scale 4, with images showing more transparent structure and improved object fidelity. In the bottom examples, representing playgrounds, the score is highest near scale 2 or 3, but they have a poor image generated. Lower scales produce blurry or semantically weak images, while higher scales enhance visual clarity but may reduce diversity. These results indicate that a moderate guidance scale, typically between 3 and 5, balances semantic alignment and image quality well. And we can also find that generation varies across different images, and tuning the guidance scale provides a simple way to control semantic accuracy and visual structure.
A.6 Limitations
We summarize key considerations and limitations of the EarthSynth-180K and EarthSynth as follows:
Limited Multispectral Generalization. This work focuses on generating optical images. However, optical images only cover the visible spectrum and lack the broader spectral information in multispectral images. This limits their use in vegetation monitoring, material classification, and environmental analysis. Although the EarthSynth framework aims for cross-satellite generalization, training and testing are done on the EarthSynth-180K dataset without standard multispectral satellites. Therefore, generalization across sensors with different spectral features is unproven. Extending the framework to handle multispectral data is an important future direction.
More Training Cost. EarthSynth’s CF-Comp strategy assumes equal training costs, but training large generative diffusion models is time- and resource-intensive. Since the CF-Comp strategy involves sample combinations in each batch, these setups may not be feasible in resource-limited or costly environments, making some comparisons more theoretical.